OpenStudy (anonymous):

earns a weekly salary and commission If she sells x $s in 1 week her earning E(x)=400+0.03x a) find E(2000) & E(15,000) what do answers represent? b)Find E(0) what it represent? c) from formula E determine what percentage she earns on good she sells

OpenStudy (ranga):

E(x)=400+0.03x To find E(2000), replace x by 2000 and compute E

OpenStudy (anonymous):

460

OpenStudy (anonymous):

so E(2000) = 460 and E(15000) = 850?

OpenStudy (ranga):

Yes, E(2000) = 460. $460 represents her weekly salary + commission if she sells $2000 worth of goods.

OpenStudy (ranga):

Yes, E(15000) = 850 is correct.

OpenStudy (anonymous):

and it also represents salary and comm?

OpenStudy (ranga):

Yes. $850 represents her weekly salary + commission if she sells $15,000 worth of goods.

OpenStudy (anonymous):

oh ok I get it and how do we find c)

OpenStudy (ranga):

First b) E(0) = 400 + (0.03)(0) = 400. $400 represents her weekly salary. There is no commission here because the goods sold is $0.

OpenStudy (anonymous):

ya I got that one

OpenStudy (ranga):

c) In the E formula x is multiplied by 0.03 where x is the dollar amount of goods sold. Therefore, 0.03 is the commission. Or the commission is 3%.

OpenStudy (anonymous):

oh ok thank you :)

OpenStudy (ranga):

you are welcome.